OpenStudy (anonymous):

Please Help! Find the value of y so that the line passing through (4, y) and (5, 3) has a slope of 1 If someone could provide a short explanation on how to do it to, I'd be greatful..

jimthompson5910 (jim_thompson5910):

m = (y2 - y1)/(x2 - x1) 1 = (y2 - y1)/(x2 - x1) 1 = (3 - y)/(5 - 4) solve for y

OpenStudy (anonymous):

i dont undertand :(

jimthompson5910 (jim_thompson5910):

I'll keep going 1 = (3 - y)/(5 - 4) 1 = (3 - y)/(1) 1*1 = 3 - y 1 = 3 - y y = ???

jimthompson5910 (jim_thompson5910):

I'm just using the slope formula, then plugging in what we're given to find the unknown value

OpenStudy (anonymous):

............

jimthompson5910 (jim_thompson5910):

where are you stuck?

OpenStudy (anonymous):

all of it...

jimthompson5910 (jim_thompson5910):

the slope formula is m = (y2 - y1)/(x2 - x1) look familiar?

OpenStudy (anonymous):

yeah, I have the formula with me right here.

jimthompson5910 (jim_thompson5910):

ok great

jimthompson5910 (jim_thompson5910):

the given slope is 1, so m = 1

jimthompson5910 (jim_thompson5910):

which is why I went from m = (y2 - y1)/(x2 - x1) to 1 = (y2 - y1)/(x2 - x1)

OpenStudy (anonymous):

okay i understand a bit now..

jimthompson5910 (jim_thompson5910):

then I plugged in the given points

jimthompson5910 (jim_thompson5910):

(x1,y1) is the first point the first point is (4,y) so you replace x1 with 4 and y1 with y ---------------------- (x2,y2) is the second point the given second point is (5,3), so you replace x2 with 5 and replace y2 with 3

jimthompson5910 (jim_thompson5910):

all of those substitutions give you 1 = (3 - y)/(5 - 4)

jimthompson5910 (jim_thompson5910):

and from there, you solve for y

OpenStudy (anonymous):

....... okay..

jimthompson5910 (jim_thompson5910):

so are things clearer?

OpenStudy (anonymous):

a bit

jimthompson5910 (jim_thompson5910):

alright keep practicing and hopefully they'll get even more clearer

OpenStudy (anonymous):

i still dont know the answer ;___;

jimthompson5910 (jim_thompson5910):

1 = (3 - y)/(5 - 4) 1 = (3 - y)/(1) 1 = 3 - y y = ____ ... fill in the blank

OpenStudy (anonymous):

i give up

jimthompson5910 (jim_thompson5910):

3 - NUMBER = 1 NUMBER = ???

OpenStudy (anonymous):

1...?

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

oh....

jimthompson5910 (jim_thompson5910):

3-y = 1 3 = 1+y 3 - 1 = y 2 = y y = 2

OpenStudy (anonymous):

i feel stupid... thank you though for tryingtoexplain it to me..

jimthompson5910 (jim_thompson5910):

no you're not stupid, so don't ever think that

jimthompson5910 (jim_thompson5910):

you just don't have enough practice

OpenStudy (anonymous):

okay, thanks

jimthompson5910 (jim_thompson5910):

yw, keep practicing